Consider the following functions. f(-2) = 5 and g(-2) = -14 Step 4 of 4: Find (-2). Answer How to enter your answer -2) =
Answer #1

Solution)

Given data is

f(-2)= 5

g(-2)= -14

But

(f/g)(x)= f(x)/g(x)

=>(f/g)(-2)= f(-2)/g(-2)

=>(f/g)(-2)= 5/14

Is the required solution .....